One type of sunburn occurs on exposure to UV light of wavelength in the vicinity of 300 nm .

a)What is the energy of a photon of this wavelength?

E =

6.62×10−19

  J  

B) What is the energy of a mole of these photons? E =

C)How many photons are in a 1.10 mJburst of this radiation? N=_______ photons
Answer #1

A)

we have below equation to be used:

Energy = Planck constant*speed of light/wavelength

=(6.626*10^-34 J.s)*(3.0*10^8 m/s)/(3*10^-7 m)

= 6.626*10^-19 J

Answer: 6.63*10^-19 J

B)

This is energy of 1 photon

Energy of 1 mol = energy of 1 photon * Avogadro's number

= 6.626*10^-19*6.022*10^23 J

= 3.99*10^5 J

= 3.99*10^2 KJ

Answer: 3.99*10^2 KJ

C)

total energy = 1.10 mJ = 0.0011m J

number of photon = total energy/energy of 1 photon

n = 0.00110/6.626*10^-19

= 1.66*10^15

Answer: 1.66*10^15 photons